OpenStudy (anonymous):

What are the intercepts of the line? x-3y=9 A. x- intercept: -3 ; y- intercept: -9 B. x- intercept: 3 ; y- intercept: 9 C. x- intercept:-9 ; y- intercept: -3 D. x- intercept: 9 , y- intercept: -3

OpenStudy (anonymous):

@Vianne

OpenStudy (vianne):

I'm sorry. I'm not familiar with this. Haha xD

OpenStudy (anonymous):

oh haha do you know anybody I can get help with maybe @Vianne

OpenStudy (a1234):

Set y to 0 for the x-intercept, and x to 0 for the y intercept.

OpenStudy (vianne):

Nope, sorry. :/

OpenStudy (anonymous):

uhh I dont understand what you mean @a1234

OpenStudy (a1234):

x-3y=9 0 - 3y = 9 Can you do that for the y-intercept?

OpenStudy (anonymous):

I know but the answer would be zero @a1234

OpenStudy (a1234):

How is that?

OpenStudy (anonymous):

0-3*0=0 @a1234

OpenStudy (a1234):

You have to set x and y to 0 separately, not at the same time. Only one variable at a time.

OpenStudy (a1234):

Can you solve -3y = 9?

OpenStudy (anonymous):

yeah

OpenStudy (a1234):

What is it?

OpenStudy (anonymous):

3

OpenStudy (er.mohd.amir):

x-3y=9 x/9 -y/3=1 then x intercept=3 and y intercept=-3 because x/a+y/b=1 x-intercept=a y-intercept=b

OpenStudy (a1234):

-3, because it is negative.

OpenStudy (a1234):

-3 is the y-intercept. Now can you do the x-intercept?

OpenStudy (anonymous):

yes thank you @a1234

OpenStudy (a1234):

YW

OpenStudy (er.mohd.amir):

sorry x-intercept is by mistake write 3.